OpenStudy (anonymous):

is this equation linear or nonlinear a + 2/b = 8

OpenStudy (raden):

nonlinear

OpenStudy (anonymous):

Thank Yu

OpenStudy (anonymous):

is ( 6,-7) a solution of 3x-4y=5

OpenStudy (nory):

3(6) - 4 (-7) = 5 18 + 28 = 5 Well, does it?